# Learning to Mask and Permute Visual Tokens for Vision Transformer Pre-Training

### v1

@@ C Additional Codebook Visualization
In Figure 4 we report a supplementary compilation of examples previously introduced in Subsection 5.3. Within this collection, distinct visual tokens successfully represent semantic concepts. Notably, the pairs of Clownfish and Lionfish, Leopard snout and Cougar snout, Yellow butterfly and Yellow flower, as well as Bicycle and Unicycle, further validate the discriminative nature of visual tokens, even for highly similar categories.
@@
In Figure 5, we also show some examples of image patches that mismatch the semantic concepts associated with the discrete tokens of the codebook. Indeed, we observe some incongruities that may be linked to the resemblance of patches to other semantic concepts. For instance, in the first row of the figure, some image patches are identified with the Peacock visual tokens while the bird species are different. The sumptuous tails of these birds can be visually associated with the characteristic features of a peacock. Similarly, in the second row of the image, the silver and gold perforations, which are associated with medieval headgear or adornments, may be perceived as comparable to the reflective surface of a carnival mask. Furthermore, within the third row of the figure, it is worth noting that several images do not precisely match the semantically correct concept of Space Shuttle. Nevertheless, these images demonstrate a semantic affinity with the broader, higher-level concept of “Space”. The images contain a variety of visual elements, including the NASA logo, astronauts, planets, and even aliens. While the specific content of these images may diverge from the targeted concept, the overall semantic theme they evoke can be seen as related to the broader concept of space exploration and travel. These observations highlight our $k$-CLIP tokenizer enables the identification and classification of complex visual concepts.
D Additional Implementation Details
Hyperparameters for Pre-Training. In In Table 7, we schematize the experimental settings adopted during the pre-training phase.
Hyperparameters for Linear Probe. Linear probing has been a widely considered proxy for assessing the effectiveness of self-supervised pre-training models. In accordance with the approach outlined in bao2022beit , we train a linear classifier head over the image-level representation output produced by the frozen pre-trained backbone. We use the class labels of the images to train the aforementioned classifier head. We train for 50 epochs using a batch size of 1024, AdamW DBLP:conf/iclr/LoshchilovH19 as optimizer, and a learning rate of $4e{-3}$ with cosine decay. The weight decay is set to $1e{-4}$. Our pre-training augmentation strategy incorporates random resizing of crops, horizontal flipping during training, and central crops during evaluation.
Computational Requirements. The pre-training experiments conducted with our MaPeT model involved the utilization of different GPU configurations. Specifically, the ViT-T, ViT-S, and ViT-B models required the deployment of 16, 32, and 64 GPUs, respectively. The pre-training process for each model took approximately one day to complete. In contrast, during the fine-tuning phase, we employed 4 GPUs for the ViT-T and ViT-S models, and 16 GPUs for the ViT-B model. The fine-tuning duration for each model was 48 hours, 36 hours, and 12 hours, respectively. For all experiments, we utilized the NVIDIA V100 GPU architecture, which was equipped with 16GB of memory and ran on IBM POWER9 AC922 CPUs operating at 3.1 GHz.
E Limitations and Societal Impact
The findings presented in Sec. A highlight the potential of MaPeT as a powerful pre-training technique for various real-world applications. The ability of MaPeT to effectively transfer knowledge across different domains opens up opportunities for more efficient and accurate visual recognition systems. By reducing the reliance on domain-specific labeled data, MaPeT offers a promising avenue for addressing the data scarcity challenge in various domains. However, it is important to acknowledge certain limitations and challenges associated with our study. MaPeT necessitates significant computational resources at training time, which can be demanding in terms of both hardware and time. These computational requirements of MaPeT can pose challenges for its widespread adoption and practical implementation, particularly in settings with limited access to high-end computing resources. While the performance benefits of MaPeT have been demonstrated in the evaluated datasets, it is important to consider the scalability and adaptability of the proposed method to even more diverse and complex domains.
Table 7: Hyperparameters for pre-training on ImageNet-1k deng2009imagenet .
@@ E Limitations and Societal Impact
Table 8: Hyperparameters for fine-tuning on ImageNet-1k deng2009imagenet .
@@
Figure 4: Visualization of image patches corresponding to the discrete tokens contained within our $k$-CLIP codebook. The codebook exhibits a high degree of semantic density and coherency, i.e., semantically similar image patches are consistently linked with the same discrete token in the visual codebook. Corresponding image patches are marked in red rectangle.
Figure 5: Visualization of image patches mismatching the semantic concept associated with the discrete tokens contained within our $k$-CLIP codebook. Corresponding image patches are marked in red rectangle.

### v2

@@ 7 Additional Codebook Visualization
In Fig. 5 we report a supplementary compilation of examples previously introduced in Sec. 5.3. Within this collection, distinct visual tokens successfully represent semantic concepts. Notably, the pairs of Clownfish and Lionfish, Leopard snout and Cougar snout, Yellow butterfly and Yellow flower, as well as Bicycle and Unicycle, further validate the discriminative nature of visual tokens, even for highly similar categories.
@@
In Fig. 4, we also show some examples of image patches that mismatch the semantic concepts associated with the discrete tokens of the codebook. Indeed, we observe some incongruities that may be linked to the resemblance of patches to other semantic concepts. For instance, in the first row of the figure, some image patches are identified with the Peacock visual tokens while the bird species are different. The sumptuous tails of these birds can be visually associated with the characteristic features of a peacock. Similarly, in the second row of the image, the silver and gold perforations, which are associated with medieval headgear or adornments, may be perceived as comparable to the surface of a carnival mask. Furthermore, within the third row of the figure, it is worth noting that several images do not precisely match the semantically correct concept of Space Shuttle. Nevertheless, these images demonstrate a semantic affinity with the broader, higher-level concept of “Space”. The images contain a variety of visual elements, including the NASA logo, astronauts, planets, and aliens. While the specific content of these images may diverge from the targeted concept, the overall semantic theme they evoke can be seen as related to the broader concept of space exploration and travel. These observations highlight our $k$-CLIP tokenizer enables the identification and classification of complex visual concepts.
@@
Table 8: Hyperparameters for pre-training on ImageNet-1k (Deng et al., 2009).
@@
Table 9: Hyperparameters for fine-tuning on ImageNet-1k (Deng et al., 2009).
Fig. 4: Visualization of image patches mismatching the semantic concept associated with the discrete tokens contained within our $k$-CLIP codebook. Corresponding image patches are marked in red rectangle.
Hyperparameter Tiny / Small / Base Size
Input resolution $512\times 512$
Peak learning rate 5e-5
Fine-tuning steps 160k
Batch size 16
Optimizer AdamW
Adam $\epsilon$ 1e-8
Adam $\beta$ (0.9, 0.999)
Layer-wise learning rate decay 0.75
Minimal learning rate 0
Warmup learning rate 5e-11
Learning rate schedule Linear
Warmup steps 1500
Weight decay 0.05
Stochastic depth 0.15
Dropout
Table 10: Hyperparameters for fine-tuning MaPeT on ADE20K (Zhou et al., 2017, 2019) for semantic segmentation tasks.
8 Additional Implementation Details
Hyperparameters for pre-training. In Table 8, we schematize the experimental settings adopted during the pre-training phase.
Hyperparameters for linear probe. Linear probing has been a widely considered proxy for assessing the effectiveness of self-supervised pre-training models. In accordance with the approach outlined in (Bao et al., 2022), we train a linear classifier head over the image-level representation output produced by the frozen pre-trained backbone. We use the class labels of the images to train the aforementioned classifier head. We train for 50 epochs using a batch size of 1024, AdamW (Loshchilov and Hutter, 2019) as optimizer, and a learning rate of $4e{-3}$ with cosine decay. The weight decay is set to $1e{-4}$. Our pre-training augmentation strategy incorporates random resizing of crops, horizontal flipping during training, and central crops during evaluation.
Hyperparameters for semantic segmentation. The hyperparameters adopted for fine-tuning MaPeT on the ADE20K (Zhou et al., 2017, 2019) dataset for semantic segmentation tasks are outlined in Table 10.
Fig. 5: Visualization of image patches corresponding to the discrete tokens contained within our $k$-CLIP codebook. The codebook exhibits a high degree of semantic density and coherency, i.e., semantically similar image patches are consistently linked with the same discrete token in the visual codebook. Corresponding image patches are marked in red rectangle.